OpenStudy (anonymous):

f(x)= {2x-3, x cannot equal 2 and 0, x=2} has a removable discontinuity. Redefine f(x) so that it is continuous at x=2.

OpenStudy (anonymous):

replace \(x\) by \(2\) and see what you get

OpenStudy (anonymous):

or better yet, just remove all the verbiage and write \[f(x)=2x-3\]

OpenStudy (anonymous):

I got 1?

OpenStudy (anonymous):

yeah ignore my first answer, just make \(f(x)=2x-3\) for all \(x\) then it is continuous for sure it is a line

OpenStudy (anonymous):

what about the 0?

OpenStudy (anonymous):

what about it?

OpenStudy (anonymous):

its the answer just f(x)=2x−3 ?

OpenStudy (anonymous):

\[f(x) = \left\{\begin{array}{rcc} 2x-3 & \text{if} & x \neq 2 \\ 0& \text{if} & x =2 \end{array} \right. \]

OpenStudy (anonymous):

the only reason this is not continuous at \(2\) is because \(2\times 2-3=-1\) and this function is defined so that \(f(2)=0\) rather than \(-1\) for some reason to get rid of that stupid discontinuity, make the function just \(f(x)=2x-3\) for all \(x\) like normal, and then it is continuous

OpenStudy (anonymous):

oh okay thank you so much:)

OpenStudy (anonymous):

can i ask you another question